OpenStudy (anonymous):

Simplify the square root of 18y to the 2nd power

OpenStudy (anonymous):

18y

OpenStudy (anonymous):

18y

OpenStudy (anonymous):

\[\sqrt{18y^2}\]

OpenStudy (anonymous):

\[y\sqrt{18}\]

OpenStudy (anonymous):

\[3y \sqrt{2}\]

OpenStudy (anonymous):

\[=3y\sqrt{2}\]